Uncertainty in position is twice the uncertainty in momentum then uncertainty in velocity will be

Grade:11

1 Answers

Arun
5 years ago
Here is a principle known as Heisenberg uncertainty principle.

according to Heisenberg uncertainty principle
= Δx . ​Δp >= h / 2
h = h/2π 
​ Δx= uncertainty in position
Δp= uncertainty in momentum= m Δv
Let , Δx =​ Δv= y
​so, y2.m=h/4π
y=[h/(4mπ)]1/2
uncertainty in momentum= y.m= [mh/4π]1/2= 1/2.[mh/π]1/2